OpenStudy (anonymous):

Is 3 part of the solution set for the inequality six times w is less than or equal to eighteen? Why or why not?

OpenStudy (anonymous):

\[6w\leq 18\iff w\leq 3\] now we check for \(w=3\) is \(3\leq 3\) true or false?

OpenStudy (anonymous):

False They are equal!

OpenStudy (anonymous):

it would be true, because the equality is included in the set

OpenStudy (radar):

I think that would read like is 3 equal OR less than 3. It would be true as 3 equals 3.